# Investigating Feature and Model Importance in Android Malware Detection: An Implemented Survey and Experimental Comparison of ML-Based Methods

### v2

@@ 1. Introduction
To address these problems, the aim of this study is to rigorously analyse the importance of feature and model choices while training ML models for Android malware detection. We primarily do this by reimplementing past works and reevaluating them using a large dataset that comprises current Android applications that we collected during the period of 2019–2021, allowing us to carry out a meaningful comparison of these methods within the current Android ecosystem and on a level playing-field. It is not feasible, nor perhaps desirable, to reimplement all the previous work in this area. Instead we focus on a group of studies which are representative of the diversity of modelling approaches used in this field, favouring those which have reported high levels of malware discrimination. This choice was guided by our previous review of published work on ML-based Android malware detection (Muzaffar et al., 2022). We also carry out a number of new experiments to fill in knowledge gaps in the existing literature, and consider the benefits of using ensemble models that combine existing approaches.
@@
  • We compile a balanced up-to-date dataset of Android malware and benign applications that is, at the time of writing this paper, the largest publicly available dataset for assessing Android anti-malware. We also share the tools  [^1] The dataset will be made publicly available once this paper is accepted. used to create this dataset to make it easier for the community to develop and assess future anti-malware.
@@
The paper is organized as follows. Section 2 describes our methodology, including dataset collection, our static and dynamic analysis tools, and our evaluation framework. This methodology is then applied to static analysis, dynamic analysis, and hybrid analysis — the three main branches of feature analysis and model building in malware detection — in Sections 3, 4 and 5, respectively. Section 7 then considers ensemble methods that combine the best models from the previous sections. Section 8 discusses the main findings of the study, and Section 9 concludes.
@@ 2. Methodology
In this section, we present out framework for assessing Android anti-malware approaches. Section 2.1 describes the dataset we use, and how it was collected. Section 2.2 describes the tool we developed for performing static and dynamic analysis of Android applications. Section 2.3 outlines the metrics we use to assess ML models. Section 2.4 motivates the choice of features and our focus on binary classification. Section 2.5 presents a core set of ML models and feature selection methods which we use throughout the study.
@@ 2.1. Dataset Collection
We collected Android applications released from 2019-2021 from the stores and repositories which are typically used by end users. The lack of availability of scripts to build datasets prompted us to build crawlers to download these applications. Through periodic use, these platform-independent Python scripts can be used to maintain an up-to-date dataset. The stores we targeted were UpToDown (Uptodown, 2024), APKMirror (APKMirror, 2024) and F-Droid (F-Droid, 2024). The scripts crawl the websites for all applications irrespective of the categories the applications are filed under. We downloaded a total of 62,000 benign applications. To the best of our knowledge, this is the most realistic and up to date benign dataset available right now.
@@ 2.2. DroidDissector
We developed an automated analysis tool, DroidDissector, to extract both static and dynamic features from Android applications. Drawing from the most prominent features highlighted in the literature, our tool attempts to provide coverage of the whole Android malware detection landscape (Muzaffar et al., 2023). The tool’s architecture is depicted in Figure 1. It comprises two sub-systems: one for extracting static features, and another for extracting dynamic features.
@@ 5. Hybrid Analysis
We begin by reimplementing a study by Kandukuru and Sharma (Kandukuru and Sharma, 2017) that combined dynamic network traffic features with static permission features. The network traffic features included all seven of the TCP features listed in Section 4.3 and the authors used a malware dataset of 1,200 applications collected from 2010 to 2011 to train a decision tree model. In our reimplementation, we trained all four core ML models. Table 28 compares their performance with Kandukuru and Sharma’s results. Since their performance is worse than models trained only on network features (see Tables tables 25 and 26), this suggests that there is no benefit to including information about static permissions.
To further investigate the use of hybrid features, we reimplemented a study based around two frequently used groups of static and dynamic features, permissions and system calls. Kapratwar et al. (Kapratwar et al., 2017) carried out experiments using these features both separately and in combination. In the case of system calls, a frequency-based representation was used. For each experiment, the authors trained RF, decision trees, Naïve Bayes, logistic regression, kNN and SVM. However, as with most past hybrid works, Kapratwar et al. used a small dataset of just 103 malware and 97 benign applications. Therefore, a reimplementation of this approach using a larger contemporary dataset is essential to understand the effect of combining two of the most prevalent features in Android malware detection.

### v3

@@ 1. Introduction
To address these problems, the aim of this study is to rigorously analyse the importance of feature and model choices while training ML models for Android malware detection. We primarily do this by reimplementing past works and reevaluating them using a large dataset that comprises current Android applications that we collected during the period of 2019–2021, allowing us to carry out a meaningful comparison of these methods within the current Android ecosystem and on a level playing-field. It is not feasible, nor perhaps desirable, to reimplement all the previous work in this area. Instead we focus on a group of studies which are representative of the diversity of modelling approaches used in this field, favouring those which have reported high levels of malware discrimination. This choice was guided by our previous review of published work on ML-based Android malware detection (Muzaffar et al., 2022). We also conduct a number of new experiments to fill in knowledge gaps in the existing literature, and consider the benefits of using ensemble models that combine existing approaches.
@@
  • We compile a balanced up-to-date dataset of Android malware and benign applications that is, at the time of writing this paper, the largest publicly available dataset for assessing Android anti-malware. We also share the tools used to create this dataset to make it easier for the community to develop and assess future anti-malware.
@@
The paper is organized as follows. Section 2 describes our methodology, including dataset collection, our static and dynamic analysis tools, and our evaluation framework. This methodology is then applied to static analysis, dynamic analysis, and hybrid analysis — the three main branches of feature analysis and model building in malware detection — in Sections 3, 4 and 5, respectively. Section 7 then considers ensemble methods that combine the best models from the previous sections. Section 8 discusses the main findings of the study, Section 9 discusses its limitations and future work, and Section 10 concludes.
@@ 2. Methodology
In this section, we present our framework for assessing Android anti-malware approaches. Section 2.1 describes the dataset that we use and how it was collected. Section 2.2 describes the tool we developed for performing static and dynamic analysis of Android applications. Section 2.3 outlines the metrics we use to assess ML models. Section 2.4 motivates the choice of features and our focus on binary classification. Section 2.5 presents a core set of ML models and feature selection methods which we use throughout the study.
@@ 2.1. Dataset Collection
We collected [^1] Dataset available at https://researchportal.hw.ac.uk/en/datasets/android-dataset-for-malware-detection Android applications released from 2019-2021 from the stores and repositories which are typically used by end users. The lack of availability of scripts to build datasets prompted us to build crawlers to download these applications. Through periodic use, these platform-independent Python scripts can be used to maintain an up-to-date dataset. The stores we targeted were UpToDown (Uptodown, 2024), APKMirror (APKMirror, 2024) and F-Droid (F-Droid, 2024). The scripts crawl the websites for all applications irrespective of the categories the applications are filed under. We downloaded a total of 62,000 benign applications. To the best of our knowledge, this is the most realistic and up to date benign dataset available right now.
@@ 2.2. DroidDissector
We developed an automated analysis tool, DroidDissector [^2] Code and scripts available at https://github.com/alisolehria/DroidDissector-A-Static-and-Dynamic-Analysis-Tool-for-Android-Malware-Detection, to extract both static and dynamic features from Android applications. Drawing from the most prominent features highlighted in the literature, our tool attempts to provide coverage of the whole Android malware detection landscape (Muzaffar et al., 2023). The tool’s architecture is depicted in Figure 1. It comprises two sub-systems: one for extracting static features, and another for extracting dynamic features.
@@ 5. Hybrid Analysis
We begin by reimplementing a study by Kandukuru and Sharma (Kandukuru and Sharma, 2017) that combined dynamic network traffic features with static permission features. The network traffic features included all seven of the TCP features listed in Section 4.3 and the authors used a malware dataset of 1,200 applications collected from 2010 to 2011 to train a decision tree model. In our reimplementation, we trained all four core ML models, using a dataset of feature vectors comprising 166 permissions and the seven TCP features. Table 28 compares their performance with Kandukuru and Sharma’s results. Since their performance is worse than models trained only on network features (see Tables tables 25 and 26), this suggests that there is no benefit to including information about static permissions.
To further investigate the use of hybrid features, we reimplemented a study based around two frequently used groups of static and dynamic features, permissions and system calls. Kapratwar et al. (Kapratwar et al., 2017) carried out experiments using these features both separately and in combination. In the case of system calls, a frequency-based representation was used. For each experiment, the authors trained RF, decision trees, Naïve Bayes, logistic regression, kNN and SVM. However, as with most past hybrid works, Kapratwar et al. used a small dataset of just 103 malware and 97 benign applications. Therefore, a reimplementation of this approach using a larger contemporary dataset is essential to understand the effect of combining two of the most prevalent features in Android malware detection. The dataset used for this reimplementation included 166 permissions and 599 system calls.
@@ 8. Discussion
9. Limitations and Future Work
The dynamic nature of the Android operating system results in new studies being published at a rapid pace. Notably, transformers are becoming increasingly popular across applied ML, and there are examples of these being applied to Android malware detection (Liu et al., 2022). However, the complexity of these models does introduce new challenges (Lones, 2021).
Our aim in this study is to sample and reimplement representative studies from the literature. It is not feasible to reimplement all existing studies, and there is inevitably some sampling bias within our approach. In particular, we focus on studies which used similar underlying features, since this enables a more meaningful basis for comparison across a standardised data set. This approach was motivated by our earlier review (Muzaffar et al., 2022) of the most common ML-based modelling approaches in Android malware detection.
It is worth noting in particular that transformers are becoming increasingly popular across applied ML, and there are examples of these being applied to Android malware detection (Oak et al., 2019; Fan et al., 2021; Saracino and Simoni, 2023; Yang et al., 2024). Unlike the approaches surveyed in this work, these approaches tend to use lower-level features, which may be constructed directly from raw data. This introduces benefits as well as challenges (Lones, 2021). An example of a transformer-based approach is the work of (Fan et al., 2021), which addresses the challenge of integrating spatial and temporal information by combining two different transformer models: one which focuses on spatial patterns and another which focuses on temporal patterns. This modelling approach also allows the system to integrate behavioural information beyond the other studies surveyed here.
There are other (non-transformer-based) studies which use broader information sources. For example, Hou et al.  (Hou et al., 2017), in their HinDroid framework, consider the relationships between applications and APIs. Similarly, Ye et al. (Ye et al., 2019) utilized features that incorporated various forms of metadata from apps and devices, including IMEI numbers.
In future work, it would be interesting to implement and compare studies that incorporate these more heterogeneous features, including recent transformer-based approaches.